OpenStudy (anonymous):

: Linear Algebra: (Diagonalization) Let matrix A = {{2,1},{-2,1}}. Find a matrix B such that B^2 = A.

OpenStudy (anonymous):

For clarification: A = \[\left[\begin{matrix}2 & 1 \\ -2 & -1\end{matrix}\right]\]

OpenStudy (phi):

you could rewrite A= S L S^-1, where L is the diagonal matrix of eigenvalues and S is the matrix of eigenvectors. Here, we see that A is singular (row 2 = - row 1), so one lambda =0. The trace 2-1= 1 is the sum of the lambdas, so the other is 1 the square root of A is S L^(1/2) S^-1, but L is has just 1 and 0 so it doesn't change. We conclude A*A= A. multiply out to show it works.

OpenStudy (anonymous):

I'm sorry, just to be sure, did you use the matrix I had in the question or in the comment? I forgot the last -1 on the matrix A in the question.

OpenStudy (phi):

I used [ 2 1] [-2 -1]

OpenStudy (anonymous):

thanks, just wanted to be sure :D

OpenStudy (anonymous):

would there be another matrix B though that if you square you get A?

OpenStudy (phi):

You know that -1 is also a square root of 1. so L^(1/2) could also be [-1 0] [ 0 0] = -L so you could deduce that B= S* -L*inv(S) = -S*L*inv(S)= -A is another solution That is, for B= [-2 -1] [ 2 1] B*B= A